# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v5

@@ Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. Combining several active learning paradigms and the core-set approach, we propose a new multi-rounds adaptation process that uses epistemic uncertainty to automate the annotation process, significantly reducing the associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples that contribute most to model uncertainty, thereby enhancing training efficiency. We define a new metric called U-WER to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
@@ 1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication [27, 17, 46]. Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents [16, 44]. This gap underscores the need for ASR systems that can handle the linguistic diversity and complexity of African languages, especially in crucial applications like healthcare. Due to the lack of representation of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names [36].
@@
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech, CommonVoices English Accented Dataset [3]), and several high-performing speech models (Wav2Vec2-XLSR-53 [10], HuBERT-Large [24], WavLM-Large [8], and NVIDIA Conformer-CTC Large (en-US) [20]). Our results show a 69.44% WER improvement while requiring on average 45% less data than established baselines. We also adapt the standard Word Error Rate (WER) to create a new metric called Uncertainty WER (U-WER) to track model adaptation to African accents.
@@
  • we show that our approach improves the WER performance by 69.44% while significantly reducing the required amount of labeled data (by $\sim$45
@@ 2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data [29, 11]. This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood [29, 28, 9, 33]. Enhancing ASR for African languages is crucial for equitable voice recognition, especially in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to these languages unique characteristics.
@@ 2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing the unlabeled examples and then selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$. The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
@@ 3.1 Datasets
We used the AfriSpeech-200 dataset [37], a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset includes 120+ African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across Africa. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US.
@@ 3.2 Methodology
4:for $k\leftarrow 1$ to $\mathcal{N}$ do
9:         $\mathcal{EUL}\leftarrow\mathcal{EUL}\cup\{(x,\text{EU}_{x})\}$
10:     end for
14:end for
[…]
3.2.1 Uncertainty WER
[…]
@@ 3.3 Experimental Design
We used various GPUs for our experiments: 6 RTX 800, and 4 A100. Training and evaluation were conducted over a month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
@@ 4 Results and Discussion
To compute the performance improvement for each domain, we take the ratio of the average of our best strategy and the baseline.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (55.90%), clinical (69.44%), and both (61.74%). Our approach also surpasses Whisper-Medium [37, 39], demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between general and clinical domains are noted, likely due to clinical samples complexity. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, aid in model learning and enhance performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the low resource availability typical of many African accents and languages.
@@ 5 Conclusion
We combined several AL paradigms, the CSA, and EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced a new metric called U-WER to track model adaptation to hard accents. Our experiments showed a remarkable 69.44% WER improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This demonstrates our approach’s efficiency and potential to significantly lower barriers to ASR technologies in underserved regions. Our method enhances model robustness and generalization across various domains, datasets, and accents, crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
@@ 6 Limitations
In discussing trade-offs (Section 4), we observed that while our approach improves performance, especially with linguistically rich accents, it is crucial for complex domains like the clinical and both domains to establish a stopping criterion that balances adaptation rounds with the selection of new data points each round. This balance is vital for refining our methodology. Regarding computational complexity, our work faced limitations due to the availability of computational resources. With better resources, we would consider implementing Deep Ensembles [30] as an alternative to our current Mc-Dropout method for estimating epistemic uncertainty.
@@ A.4 Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with both random and EU-Most selection strategies.

### v6

@@ Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. Combining several active learning paradigms and the core-set approach, we propose a new multi-rounds adaptation process that uses epistemic uncertainty to automate the annotation process, significantly reducing the associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples contributing most to model uncertainty, enhancing training efficiency. We define a new U-WER metric to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 27% WER relative average improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
@@ 1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication [27, 17, 46]. Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents [16, 44]. This gap underscores the need for ASR systems that can handle the linguistic diversity and complexity of African languages, especially in crucial applications like healthcare. Due to the lack of representations of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names [36].
@@
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech, CommonVoices English Accented Dataset [3]), and several high-performing speech models (Wav2Vec2-XLSR-53 [10], HuBERT-Large [24], WavLM-Large [8], and NVIDIA Conformer-CTC Large (en-US) [20]). Our results show a 27% Word Error Rate (WER) relative average improvement while requiring on average 45% less data than established baselines. We also adapt the standard WER to create a new metric called Uncertainty WER (U-WER) to track model adaptation to African accents.
@@
  • we show that our approach improves the relative average WER performance by 27% while significantly reducing the required amount of labeled data (by $\sim$45%),
@@ 2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data [29, 11]. This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood [29, 28, 9, 33]. Enhancing ASR for African languages is crucial for equitable voice recognition, especially in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to the unique characteristics of these languages.
@@ 2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing the unlabeled examples and selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$ ($W$ are the parameters of the model). The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
@@ 3.1 Datasets
We used the AfriSpeech-200 dataset [37], a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset includes 120+ African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing African regional diversity. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US (see Table 1).
@@ 3.2 Methodology
[…]
3.2.1 Uncertainty WER
[…]
4:for $k\leftarrow 1$ to $\mathcal{N}$ do
9:         $\mathcal{EUL}\leftarrow\mathcal{EUL}\cup\{(x,\text{EU}_{x})\}$
10:     end for
14:end for
@@ 3.3 Experimental Design
For our experiments, we used 6 RTX8000 GPUs and 4 A100 GPUs. Training and evaluation were conducted over a month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
@@ 4 Results and Discussion
To assess the performance improvement for each domain, we compute the relative average improvement $\text{RIA}_{wer,d}=\left(\frac{b_{wer}^{d}-s_{wer}^{d}}{b_{wer}^{d}}\right)\times 100\%$ where $b_{wer}^{d}$ and $s_{wer}^{d}$ are the average WER respectively of the baseline, and the best selection strategy, in a domain $d\in\{general,clinical,both\}$. A higher percentage reflects a higher improvement of our approach.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (27.00%), clinical (15.51%), and both (26.56%). Our approach also surpasses Whisper-Medium [37, 39], demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between general and clinical domains are noted, likely due to clinical samples complexity. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, aid in model learning and enhance performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the low-resource availability typical of many African accents and languages.
@@ 5 Conclusion
We combined several AL paradigms, the CSA, and EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced the U-WER metric to track model adaptation to hard accents. Our experiments showed a remarkable 27% WER ratio improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This demonstrates our approach’s efficiency and potential to lower the barriers to ASR technologies in underserved regions significantly. Our method enhances model robustness and generalization across various domains, datasets, and accents, which are crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
@@ 6 Limitations
In discussing trade-offs (Section 4), we noted that while our approach enhances performance, particularly with linguistically rich accents, a stopping criterion is essential for complex domains like the clinical one, to balance adaptation rounds with the pool size. With better resources, we would consider implementing Deep Ensembles [30] as an alternative to our current MC-Dropout method for estimating epistemic uncertainty, and leverage other acquisition functions (such as BALD, BatchBALD) highlighted in this work.
@@ A.4 Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with random and EU-Most selection strategies.